Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

  

  Sales $ 22,400    
  Variable expenses 12,800    
  Contribution margin 9,600    
  Fixed expenses 7,968    
  Net operating income $ 1,632

What is the break-even point in unit sales?
Answer #1

Contribution Margin per unit = 9600 / 1000 = 9.60

Break even units

= Fixed cost / Contribution margin per unit

= 7968/9.60

= 830 units